A lightbulb radiates most strongly at a wavelength of about 2940 nanometers. How hot is its filament?
Answer #1

According to Wein’s law,

2940 nanometers = 3*10^6nm-K/T

Thus, T = 3*10^6nm-K* 2940 nm = 3000/2940 K = 1020.4 K--> The temperature of the filament